# BranchGRPO: Stable and Efficient GRPO with Structured Branching in Diffusion Models

### v3

@@ 4.4 Ablation Studies
Figure 5(a) shows the effect of varying the branch correlation $s$. Smaller values ($s=1.0,2.0$) limit exploration and lead to slower reward growth, while very large values ($s=8.0$) destabilize early training. A moderate setting ($s=4.0$) achieves the best trade-off, reaching the highest reward and stable convergence, confirming that stochastic branching is necessary but should be carefully tuned.
Branching Steps.
@@
We next vary the positions of split timesteps (Figure 5(b)). Early splits such as $(0,3,6,9)$ promote faster reward increase in the early stage, whereas later splits like $(9,12,15,18)$ delay exploration and yield lower rewards. Intermediate schedules (e.g., $(3,6,9,12)$) balance efficiency and reward quality, suggesting that early splits are generally more effective for exploration.
Branch Density.
@@
Finally, we compare different densities of split points while keeping the overall horizon fixed (Figure 5(c)). Although all configurations eventually converge to similar reward levels, denser splits (e.g., $(0,3,6,9)$) accelerate early training, while sparser configurations converge more slowly. This indicates that increasing the density of branching in the early phase improves sample efficiency without harming stability.
Figure 5: Ablation on branching hyperparameters without pruning. (a) Moderate correlation (s=4s{=}4) best balances exploration and stability. (b) Early split schedules yield higher rewards than late splits. (c) Denser early branching accelerates reward growth while converging to similar final levels.
Figure 6: Ablations on pruning and reward fusion. Left: pruning policies (none, depth, width-Parent-Top1, width-Extreme-$b$). Right: reward-fusion strategies (uniform vs. path-weighted, with/without width pruning). Depth pruning achieves the best final reward, while path-weighted fusion improves stability across settings.
Path-weighted vs. Uniform Fusion.
@@
Figure 6 (right) compares uniform averaging with path-probability weighting under identical training settings. Uniform averaging shows higher variance and a clear late-stage plateau, whereas path-weighted fusion delivers consistently higher and more stable rewards throughout training. This empirically supports Sec. 3.3: uniform averaging is exploration-friendly but noisy, while path weighting aligns credit assignment with the behavior distribution and improves convergence at no extra cost.
Pruning.
@@
Figure 6 (left) evaluates pruning applied after depth-wise normalization and only for backprop. Depth pruning (sliding window over denoising steps) achieves the best final reward, consistently surpassing the no-pruning baseline by reallocating compute to informative early steps while preserving forward sampling diversity. Width pruning (Parent-Top1) halves gradient updates and yields the smoothest, lowest-variance curve, but its final reward is slightly below depth pruning. Width pruning (Extreme-$b$) keeps the globally best and worst $b$ leaves; it injects strong positive/negative signals and remains competitive at the end, but exhibits higher variance than the other settings.
@@ 4.5 Video Generation Results
We further evaluate BranchGRPO on video generation using WanX-1.3B I2V (Wan et al., 2025). We compare three settings: the base model, DanceGRPO, and BranchGRPO. Figure 7 presents qualitative frames and quantitative reward curves.
Qualitatively, the base model exhibits severe temporal flickering, while DanceGRPO partially improves consistency but still struggles with detail preservation. BranchGRPO produces sharper frames with coherent objects across time.
Quantitatively, BranchGRPO achieves faster convergence and higher rewards, measured by the Video-Align motion quality metric (Liu et al., 2025b), confirming the benefit of branching rollouts for video generation.
Figure 7: Video generation results on WanX-1.3B I2V. Left: qualitative frame comparisons across three settings. (a) Without GRPO: severe temporal flickering and deformation. (b) DanceGRPO: improved but still inconsistent and blurry details. (c) BranchGRPO: sharper and temporally coherent frames. Right: reward curves showing faster convergence and higher final rewards with BranchGRPO compared to DanceGRPO.
@@ 5 Conclusion
We introduced BranchGRPO, a reinforcement learning framework that mitigates reward sparsity and rollout inefficiency in diffusion and flow models. By reformulating rollouts as tree-structured trajectories with reward fusion and lightweight pruning, BranchGRPO achieves denser credit assignment, faster convergence, and higher alignment quality at lower cost. These results establish BranchGRPO as a stable, efficient alternative to existing GRPO variants and a step toward scalable preference optimization for complex generative tasks.
@@ Appendix A Appendix
  • Section D: Additional experiments, including ablations, more text-to-image and image-to-video results.
@@ Appendix B Hyperparameter Settings
Appendix C Theoretical Analysis (Detailed Proofs)
@@ C.1 Branch Noise Construction and Boundary Distribution Preservation
At a split step $i\in B$, we construct $K$ branch noises as
$\xi_{b}=\frac{\xi_{0}+s\,\delta\xi_{b}}{\sqrt{1+s^{2}}},\qquad\xi_{0},\delta\xi_{b}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}(0,I),$ (7)
so that $\xi_{b}\sim\mathcal{N}(0,I)$ for each $b$, and $\mathrm{Cov}(\xi_{b},\xi_{b^{\prime}})=\tfrac{1}{1+s^{2}}I$ for $b\neq b^{\prime}$.
@@
Conditioned on prefix $(z_{0},\eta_{0},\dots,\eta_{i-1})$, each branch $z^{(b)}_{N}$ generated by the split rule has the same distribution as a baseline SDE sample $z_{N}$.
@@
Under the branching construction above, for any set of split steps $B$, each leaf $z^{(b)}_{N}$ has the same marginal law as a baseline SDE rollout. Hence, branching does not alter the final generator distribution.
@@ C.4 Continuous Rewards and Concentration
Appendix D Additional Experiments
D.1 More quantitative result
@@
D.2 More Text2Image Results
@@
D.3 More Image2Video Results

### v4

@@ 4.4 Ablation Studies
Branch Correlation. Figure 5(a) shows the effect of varying the branch correlation $s$. Smaller values ($s=1.0,2.0$) limit exploration and lead to slower reward growth, while very large values ($s=8.0$) destabilize early training. A moderate setting ($s=4.0$) achieves the best trade-off, reaching the highest reward and stable convergence, confirming that stochastic branching is necessary but should be carefully tuned.
@@
Branching Steps. We next vary the positions of split timesteps (Figure 5(b)). Early splits such as $(0,3,6,9)$ promote faster reward increase in the early stage, whereas later splits like $(9,12,15,18)$ delay exploration and yield lower rewards. Intermediate schedules (e.g., $(3,6,9,12)$) balance efficiency and reward quality, suggesting that early splits are generally more effective for exploration.
@@
Branch Density. Finally, we compare different densities of split points while keeping the overall horizon fixed (Figure 5(c)). Although all configurations eventually converge to similar reward levels, denser splits (e.g., $(0,3,6,9)$) accelerate early training, while sparser configurations converge more slowly. This indicates that increasing the density of branching in the early phase improves sample efficiency without harming stability.
Figure 5: Ablation studies of BranchGRPO. Moderate branch correlation, early and denser splits improve reward growth; path-weighted fusion enhances stability; depth pruning achieves the best final reward; and the hybrid ODE–SDE schedule provides the fastest training speed while remaining stable.
@@
Reward Fusion Strategies. Figure 5(d) compares uniform averaging with path-probability weighting under identical training settings. Uniform averaging shows higher variance and a clear late-stage plateau, whereas path-weighted fusion delivers consistently higher and more stable rewards throughout training. This empirically supports Sec. 3.3: uniform averaging is exploration-friendly but noisy, while path weighting aligns credit assignment with the behavior distribution and improves convergence at no extra cost.
@@
Pruning Strategies. Figure 5(e) evaluates pruning applied after depth-wise normalization and only for backprop. Depth pruning (sliding window over denoising steps) achieves the best final reward, consistently surpassing the no-pruning baseline by reallocating compute to informative early steps while preserving forward sampling diversity. Width pruning (Parent-Top1) halves gradient updates and yields the smoothest, lowest-variance curve, but its final reward is slightly below depth pruning. Width pruning (Extreme-$b$) keeps the globally best and worst $b$ leaves; it injects strong positive/negative signals and remains competitive at the end, but exhibits higher variance than the other settings.
Hybrid ODE–SDE. To further explore depth pruning, inspired by MixGRPO (Li et al., 2025a), we design a hybrid ODE–SDE schedule: all branching steps are preserved as SDE, while a sliding window determines additional SDE steps, with the remaining updates replaced by ODE. Figure 5(f) shows that this scheme achieves the fastest speedups (148s vs. 289s for MixGRPO vs. 469s for DanceGRPO ) while maintaining stable and fast reward growth.
4.5 Scaling with Branch
DanceGRPO scales poorly: generating 81 rollout samples requires over 2400s per iteration, whereas BranchGRPO achieves the same scale in only 680s, making large-scale scaling feasible. We investigate two settings: scaling the branch factor (K=2,3,4K{=}2,3,4 yielding 16, 81, and 256 leaves) and scaling the number of branching steps ($3,4,5$ splits yielding 8, 16, and 32 leaves under K=2K{=}2).
As shown in Figure 6, scaling along both dimensions leads to clear and substantial gains in reward growth and final performance. Larger branch factors and more branching steps consistently push the reward curves higher, with improvements becoming increasingly pronounced as the rollout tree expands. This demonstrates that BranchGRPO can effectively generate additional samples, making large-scale scaling both practical and beneficial without compromising stability.
Figure 6: Impact of scaling branch rollouts in BranchGRPO. Larger branch factors (a) and more branching steps (b) consistently improve reward, while iteration time (c) increases with leaves.
Figure 7: Video generation results on WanX-1.3B. Left: qualitative frame comparisons across three settings. Right: reward curves showing faster convergence and higher final rewards with BranchGRPO compared to DanceGRPO.
@@ 4.6 Video Generation Results
We further evaluate BranchGRPO on video generation using WanX-1.3B (Wan et al., 2025), with the Video-Align motion quality (Liu et al., 2025b) as reward.
As shown in Figure 7, the base model exhibits severe temporal flickering and deformation, while DanceGRPO improves consistency but still produces blurry details. BranchGRPO generates sharper and more coherent frames across time, and the reward curves demonstrate faster convergence and higher final rewards compared to DanceGRPO.
These results highlight that branching rollouts are particularly effective for video generation, where reward sparsity and temporal coherence are especially challenging. In practice, BranchGRPO also improves efficiency: each iteration takes only about 8 minutes, compared to 20 minutes for DanceGRPO. Additional visual examples are provided in the supplementary material.
@@ 5 Conclusion
We presented BranchGRPO, a tree-structured reinforcement learning framework that overcomes inefficiency and sparse credit assignment in GRPO by introducing branching rollouts with shared prefixes, reward fusion with depth-wise advantages, and lightweight pruning. Experiments on image and video generation show that BranchGRPO delivers faster, more stable, and higher-quality alignment, establishing structured branching as a practical and scalable path for RLHF in diffusion and flow models.
@@ Appendix A Appendix
  • Section E: Additional experiments, including ablations, more text-to-image and image-to-video results.
@@ Appendix B Hyperparameter Settings
Appendix C Theoretical Analysis
@@ C.1 Branch Noise Construction and Boundary Distribution Preservation
At a split step $i\in\mathcal{B}$, we construct $K$ branch noises as
$\xi_{b}=\frac{\xi_{0}+s\,\eta_{b}}{\sqrt{1+s^{2}}},\qquad\xi_{0},\eta_{b}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}(0,I),$ (7)
with fresh $(\xi_{0},\{\eta_{b}\})$ drawn at each split step and no cross-time sharing. Then $\xi_{b}\sim\mathcal{N}(0,I)$ for each $b$, and $\mathrm{Cov}(\xi_{b},\xi_{b^{\prime}})=\tfrac{1}{1+s^{2}}I$ for $b\neq b^{\prime}$.
@@
Assuming independent noises across time steps and no cross-time reuse of the shared component, conditioned on prefix $(z_{0},\eta_{0},\dots,\eta_{i-1})$, each branch $z^{(b)}_{N}$ generated by the split rule has the same distribution as a baseline SDE sample $z_{N}$.
@@
Under the branching construction above, for any set of split steps $\mathcal{B}$, each leaf $z^{(b)}_{N}$ has the same marginal law as a baseline SDE rollout. Hence, branching does not alter the final generator distribution.
@@ C.2 Reward Fusion: Unbiasedness and Variance Reduction
Remark. In practice (Eq. (4) in the main text), we adopt a softmax weighting $w_{\ell}\propto\exp(\beta s_{\ell})$ based on path log-probabilities $s_{\ell}$. This can be interpreted as a temperature-smoothed variant of SNIS, where $\beta$ controls the sharpness of importance weights. Although no longer strictly unbiased, this form provides stable training and reduces the influence of low-probability noisy leaves.
@@ C.4 Continuous Rewards and Concentration
Appendix D Experiment Setup Details
This section provides extended details of our experimental setup, complementing the summary in Section 4.3 of the main paper. We describe dataset splits, model backbones, branching configurations, pruning strategies, and evaluation protocols.
Datasets.
We conduct experiments on HPDv2.1 (Wu et al., 2023), which contains 103k training prompts and 400 balanced test prompts across diverse styles. For video generation, we adopt WanX-1.3B (Wan et al., 2025) on the image-to-video (I2V) task, following the standard evaluation with Video-Align (Liu et al., 2025b).
Model Backbone.
We use FLUX.1-Dev (batifol2025flux) as the text-to-image backbone. All experiments are initialized from the same pretrained checkpoint. For video tasks, we build on the WanX-1.3B model.
Branching Configurations.
We set default tree depth to $d=20$ and branching factor $K=2$, yielding 16 leaves per rollout before pruning. Branching steps $\mathcal{B}$ follow three schedules: Dense $(0,3,6,9)$, Mixed $(0,4,8,12)$, and Sparse $(0,5,10,15)$. We sweep correlation $s\in\{0,1,2,4,8\}$ to study its effect. All SDE steps are retained in forward rollouts; pruning is only applied after the final branching depth, while width pruning selects among leaves at the last split. For fair comparison, depth and width pruning drop the same number of gradient steps.
Evaluation Metrics.
We report efficiency metrics including per-iteration wall-clock time, effective number of function evaluations (NFE), and GPU-hours. For quality, we use HPS-v2.1 (Wu et al., 2023), PickScore (Kirstain et al., 2023), and ImageReward (Xu et al., 2023). For video generation, we use Video-Align (Liu et al., 2025b). Additional metrics (KID (Bińkowski et al., 2018), MMD2 (Gretton et al., 2012), IntraFID (kynkaanniemi2019improved), CLIP-based metrics (radford2021clip)) are reported in the diversity analysis.
Appendix E Additional Experiments
E.1 More quantitative result
@@
E.2 More Text2Image Results
@@
E.3 More Image2Video Results
@@ E.4 Failure Cases
Appendix F Discussion, Limitations, and Future Work
While our results demonstrate the stability and efficiency benefits of BranchGRPO, several open directions remain.
Discussion.
BranchGRPO introduces structured branching and pruning into GRPO training, which we have shown to improve both efficiency and alignment. Our ablations suggest that the choice of branching schedule and pruning strategy can substantially affect reward stability, highlighting the importance of principled tree design. Moreover, reward fusion provides stable gradients in practice, but its bias–variance tradeoff under different weighting schemes warrants further theoretical analysis.
Limitations.
First, our study is limited to image alignment on HPDv2.1 and video alignment on WanX-1.3B. Although these benchmarks are representative, broader validation on more diverse datasets and larger-scale video models is necessary. Second, our current design relies on pre-specified branching schedules and fixed correlation strengths; while effective, this may underexplore adaptive structures that respond to sample difficulty or reward uncertainty. Finally, our pruning strategies reduce gradient cost but still require full forward rollouts, leaving memory and bandwidth overhead partially unaddressed.
Future Work.
Several promising directions extend beyond the present scope. (1) Dynamic branching. Instead of fixed hyperparameters, one can design adaptive policies that adjust branch factor, correlation, or pruning windows on-the-fly based on sample difficulty or intermediate rewards, enabling more efficient rollouts. (2) Beyond diffusion models. The branching framework could naturally transfer to other generative paradigms, including diffusion-based LLMs and multimodal foundation models. (3) Scaling to long-horizon video. While initial experiments on WanX-1.3B I2V show benefits, more extensive validation on high-resolution, long-duration video generation tasks is required. (4) Robotics and action generation. Tree-structured rollouts are naturally suited to robotics, where intermediate states provide dense and verifiable rewards (e.g., task success signals). Extending BranchGRPO to robotic action generation and embodied video generation learning could open a promising direction for embodied AI.